OpenStudy (anonymous):

solve 2^(3x-4) = 5 x=____?

OpenStudy (solomonzelman):

use LOG .

OpenStudy (anonymous):

how would i set it up?

OpenStudy (anonymous):

yes, apply log and then you can kick that exponent to the front of the log

OpenStudy (anonymous):

write log in front of both terms have you used logarithms before?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

log 2^(3x-4) = log 5

OpenStudy (anonymous):

yes, and now we can move the exponent to the front of the log

OpenStudy (anonymous):

3x-4 log 2 = log 5

OpenStudy (anonymous):

remember your brackets though

OpenStudy (solomonzelman):

yes, andrew, good so far.

OpenStudy (anonymous):

and now it shouldn't be so confusing... you just need to do some algebra and solve for x

OpenStudy (anonymous):

(3x-4) log(2) = log(5)?

OpenStudy (anonymous):

great

OpenStudy (solomonzelman):

yes, it is good....

OpenStudy (solomonzelman):

1) divide both sides by log(2)

OpenStudy (anonymous):

would the equation be ...

OpenStudy (anonymous):